OpenStudy (anonymous):

How much would $200 invested at 6% interest compounded annually be worth after 5 years? Round your answer to the nearest cent. Do not include units in your answer.

OpenStudy (anonymous):

Total = P(1+r)^n P = principal = $200 r = annual interest rate = 0.06 n = number of years = 5 Total = 200(1+0.06)^5 Plug all the numbers in and calculate it with a calculator.